# RiskAtlas: Exposing Domain-Specific Risks in LLMs through Knowledge-Graph-Guided Harmful Prompt Generation

### v1

@@ Abstract
Large language models (LLMs) are increasingly applied in specialized domains such as finance and healthcare, where they introduce unique safety risks. Domain-specific datasets of harmful prompts remain scarce and still largely rely on manual construction; public datasets mainly focus on explicit harmful prompts, which modern LLM defenses can often detect and refuse. In contrast, implicit harmful prompts—expressed through indirect domain knowledge—are harder to detect and better reflect real-world threats. We identify two challenges: transforming domain knowledge into actionable constraints and increasing the implicitness of generated harmful prompts. To address them, we propose an end-to-end framework that first performs knowledge-graph-guided harmful prompt generation to systematically produce domain-relevant prompts, and then applies dual-path obfuscation rewriting to convert explicit harmful prompts into implicit variants via direct and context-enhanced rewriting. This framework yields high-quality datasets combining strong domain relevance with implicitness, enabling more realistic red-teaming and advancing LLM safety research. We release our code and datasets at GitHub. [^1] https://github.com/huawzheng/RiskAtlas
RiskAtlas: Exposing Domain-Specific Risks in LLMs through Knowledge-Graph-Guided Harmful Prompt Generation
@@ 2.2 Jailbreak and Obfuscation Methods
Figure 1: RiskAtlas: An end-to-end synthesis framework for domain-specific harmful prompt generation.
@@ 3 Methodology
Figure 1 illustrates RiskAtlas, an end-to-end pipeline for domain-specific harmful prompt synthesis. A domain knowledge graph is built from Wikidata with root selection and scale control for coverage. Guided by retrieved entities and few-shot exemplars, we generate explicit prompts, filter for toxicity and fluency, then apply dual-path obfuscation (direct and context-card rewriting) to yield stealthier, domain-relevant attacks.
@@ 4 Experiments
Model AdvBench Do-Not-Answer HarmfulQA RA-Origin $RA-Implicit\checkmark$
@@
Table 1: Evaluation of attack success rate (ASR, %) on public benchmarks and our RiskAtlas (RA).
Metric AdvBench Do-Not-Answer HarmfulQA RA-Origin $RA-Implicit\checkmark$
@@ 4.1 Experimental Setup
w/o SFT AdvBench Do-Not-Answer RA-Origin $RA-Implicit\checkmark$
@@
Metric w/o SFT AdvBench Do-Not-Answer RA-Origin $RA-Implicit\checkmark$
@@ 4.3 Performance Comparison on Safety Fine-Tuning
Explicit attack performance. We evaluate models on general-domain harmful prompts (e.g., HarmfulQA, CatQA-en) to assess whether domain-specific data degrades alignment. As shown in the upper part of Table 3, RiskAtlas performs on par with or better than public datasets under explicit attacks. For example, on HarmfulQA, RA-Origin achieves 9.0% ASR, compared to 11.0% for AdvBench and 12.5% for Do-Not-Answer; on CatQA-en, RA-Origin attains 7.0% ASR, matching AdvBench and improving over Do-Not-Answer (12.0%). These results indicate that alignment under domain specialization does not compromise robustness to explicit harmful prompts.
@@ 5 Conclusion
We present a scalable pipeline that integrates knowledge-graph–guided generation with dual-path obfuscation rewriting to build domain-specific harmful-prompt datasets. By grounding synthesis in structured domain knowledge, RiskAtlas systematically surfaces high-risk entities and extends coverage beyond surface-level vulnerabilities. The obfuscation stage transforms explicit queries into realistic, stealthy variants that better reflect real-world misuse. Extensive experiments in medicine, finance, law, and education show that RiskAtlas outperforms existing benchmarks and generalizes across models and domains.

### v2

# StealthGraph: Exposing Domain-Specific Risks in LLMs through Knowledge-Graph-Guided Harmful Prompt Generation

@@ Abstract
Large language models (LLMs) are increasingly applied in specialized domains such as finance and healthcare, where they introduce unique safety risks. Domain-specific datasets of harmful prompts remain scarce and still largely rely on manual construction; public datasets mainly focus on explicit harmful prompts, which modern LLM defenses can often detect and refuse. In contrast, implicit harmful prompts—expressed through indirect domain knowledge—are harder to detect and better reflect real-world threats. We identify two challenges: transforming domain knowledge into actionable constraints and increasing the implicitness of generated harmful prompts. To address them, we propose an end-to-end framework that first performs knowledge-graph-guided harmful prompt generation to systematically produce domain-relevant prompts, and then applies dual-path obfuscation rewriting to convert explicit harmful prompts into implicit variants via direct and context-enhanced rewriting. This framework yields high-quality datasets combining strong domain relevance with implicitness, enabling more realistic red-teaming and advancing LLM safety research. We release our code and datasets at GitHub. [^1] https://github.com/huawzheng/StealthGraph
StealthGraph: Exposing Domain-Specific Risks in LLMs through Knowledge-Graph-Guided Harmful Prompt Generation
@@ 2.2 Jailbreak and Obfuscation Methods
Figure 1: StealthGraph: An end-to-end synthesis framework for domain-specific harmful prompt generation.
@@ 3 Methodology
Figure 1 illustrates StealthGraph, an end-to-end pipeline for domain-specific harmful prompt synthesis. A domain knowledge graph is built from Wikidata with root selection and scale control for coverage. Guided by retrieved entities and few-shot exemplars, we generate explicit prompts, filter for toxicity and fluency, then apply dual-path obfuscation (direct and context-card rewriting) to yield stealthier, domain-relevant attacks.
@@ 4 Experiments
Model AdvBench Do-Not-Answer HarmfulQA SG-Origin $SG-Implicit\checkmark$
@@
Table 1: Evaluation of attack success rate (ASR, %) on public benchmarks and our StealthGraph (SG).
Metric AdvBench Do-Not-Answer HarmfulQA SG-Origin $SG-Implicit\checkmark$
@@ 4.1 Experimental Setup
w/o SFT AdvBench Do-Not-Answer SG-Origin $SG-Implicit\checkmark$
@@
Metric w/o SFT AdvBench Do-Not-Answer SG-Origin $SG-Implicit\checkmark$
@@ 4.3 Performance Comparison on Safety Fine-Tuning
Explicit attack performance. We evaluate models on general-domain harmful prompts (e.g., HarmfulQA, CatQA-en) to assess whether domain-specific data degrades alignment. As shown in the upper part of Table 3, StealthGraph performs on par with or better than public datasets under explicit attacks. For example, on HarmfulQA, SG-Origin achieves 9.0% ASR, compared to 11.0% for AdvBench and 12.5% for Do-Not-Answer; on CatQA-en, SG-Origin attains 7.0% ASR, matching AdvBench and improving over Do-Not-Answer (12.0%). These results indicate that alignment under domain specialization does not compromise robustness to explicit harmful prompts.
@@ 5 Conclusion
We present a scalable pipeline that integrates knowledge-graph–guided generation with dual-path obfuscation rewriting to build domain-specific harmful-prompt datasets. By grounding synthesis in structured domain knowledge, StealthGraph systematically surfaces high-risk entities and extends coverage beyond surface-level vulnerabilities. The obfuscation stage transforms explicit queries into realistic, stealthy variants that better reflect real-world misuse. Extensive experiments in medicine, finance, law, and education show that StealthGraph outperforms existing benchmarks and generalizes across models and domains.